Question One.
On January 1, 2016, ZANACO Bank was incorporated with the authorized capital
of K3, 000, 000, K2 Par Value common stock and 200,000 K100 par value 10%
Cumulative preferred stock. The following transactions relating to the stockholders
equity occurred in the first year of operations.
• On January 1st 2016, the board authorized the issuance of K1, 000, 000
common stock shares of these 500, 000 shares were issued for cash at K3 per
share, while 500, 000 were subscribed for K4 per share.
• On January 2nd 2016, a further 1 million common stock shares were issued in
exchange for the office building valued at k3,500,000
• On January 3rd 2016the company issued 100,000 10% cumulative preference
shares at k120 per share
• On January 30th U.B.A received full payment for the shares subscribed for on
January 1st.
Required;
a. Prepare journal entries to record the above transactions. ( 5 marks)
b. Prepare the stockholders equity section of ZANACO statement of Financial
Position as at December 2016.
c. ZANACO made losses in the first two years (2016 and 2017) of operation but
in the third year 2018 it made a profit of K50 Million. Show how this profit
will be shared among the shareholders if the board decides that 25% of the
profits should be retained. (2 Marks)
d. Based on the information given in “C” above, calculate the earnings per share
“EPS” for common shareholders (1 mark)
Answer #1

1 Jan 2016

(i) Bank A/c ...... Dr. K 3500000 Working (500000* 3) + (500000* 4) = K 3500000

To Common stock K 3500000

( Being common stock shares Issued)

2 Jan 2016 Building A/c ...... Dr. K 3500000

To common stock K 3500000

( Being Building Purchased and shares issued)

3 Jan 2016 Bank A/c ...... Dr. K 12,000,000

To 10 % Preference share A/c K 12,000,000

( Being Preference share issued)

(B) Statement of holding

Common stock = K 4000000

Preference share =   K 10000000

Premium = K 1500000 + K 1500000 + K 2000000 = K 5000000

(C) Now amount to be Divided

Earnings available for common stock holders

Earnings - Distributed to preference share holders

K 500000000 - K 1000000 = 49000000

Earnings to be retained

25% of K 49000000 = K 12250000

Earnings to be distributed = K 36750000

( D ) EPS = K 36750000 / 2000000 = 18.375